# Towards User-Focused Research in Training Data Attribution for Human-Centered Explainable AI

### v1

@@ 5. Discussion
This work presents a needfinding study with AI practitioners for training data attribution (TDA) explanations to inform user-focused TDA research in a top-down approach. We summarise our findings to provide answers to the initial research questions (§ 5.1), identify research gaps that need to be addressed for more user-focused TDA (§ 5.2) and discuss limitations (§ 5.3).
@@ 5.1. Answers to the research questions
5.1.1. RQ1: What do users need TDA explanations for?
To answer this question, we conducted the interview study (§ 4.1). Our findings suggest that TDA as an explanation method has the potential to serve similar high-level purposes as other explanation techniques, namely enabling users to understand model behaviour, especially unexpected model behaviour. By offering insights into model behaviour, TDA can assist end-users in calibrating trust by providing context for model decisions and support model developers in debugging errors by identifying root causes in the training data. This data-centric approach is particularly valuable in model development, leading us to identify model developers as the primary audience for TDA explanations and model debugging as the key use case.
5.1.2. RQ2: What do users need from TDA explanations?
To answer this question, we designed a survey study that extracts user preferences for the information represented by TDA (§ 4.4). Our study reveals that users need different types of information to understand model behaviour, depending on their hypotheses of why the model errs during development. This necessitates TDA explanations that are flexible and adaptable to individual needs. These needs vary based on the application domain and the user’s level of expertise. Experienced users, may form hypotheses about model errors based on their past experiences. In contrast, less experienced users may form hypotheses based on what they learnt previously about machine learning models. Users have a preference for group attribution, as groups provide more information than individual samples. Additionally, users need TDA explanations to be reliable considering how data-centric actions might affect the model as a whole. Reliability entails two main aspects: First, the information needs to be reliable for it to be actionable and effective for the task (i.e., model debugging). Second, TDA explanations should also account for any broader changes in the model’s behaviour that may affect the attribution to training samples.
5.1.3. RQ3: To what extent do existing TDA approaches align with user needs?
To answer this question, we reflect on existing TDA approaches in light of the identified needs: TDA explanations need to be adaptable to the user and provide reliable and actionable insight while maintaining consistency with the overall model behaviour. From these needs, we identify different TDA tasks that users need, next to the overarching need for reliability (see Table 5).
@@ 5.1.3. RQ3: To what extent do existing TDA approaches align with user needs?
| Influence attribution (e.g., (Koh and Liang, 2017; Pruthi et al., 2020; Guo et al., 2021; Schioppa et al., 2022; Grosse et al., 2023; Park et al., 2023)) | Removal | Sample loss | 1 | 1 | 0 |
@@
Traditional TDA measures the influence of a training sample on model predictions by analysing the impact of its removal through leave-one-out retraining (Hampel, 1974). This method essentially evaluates model predictions in the context of a counterfactual change in the training data. It has been studied across different tasks, such as identifying mislabelled data, understanding adversarial vulnerabilities, and detecting domain mismatches (Koh and Liang, 2017). Mislabel identification, in particular, is a common evaluation strategy employed in several studies (Pruthi et al., 2020; Park et al., 2023; Schioppa et al., 2022). Consequently, traditional TDA partially addresses the need for adaptability to different use cases. However, since TDA quantifies the relevance of a training sample by its removal, it is particularly suited for scenarios where data elimination is necessary, such as malicious samples. It remains an open question whether alternative actions, such as changing labels or upsampling a data region, might provide improved explanations in terms of actionability. Our study indicates that users may have diverse ideas regarding the root causes of errors, and traditional TDA only addresses one of these perspectives.
@@
Regarding the need for reliable, actionable, and effective information, existing TDA methods have been criticised for their low reliability in deep learning (Nguyen et al., 2023; Basu et al., 2021; Bae et al., 2022). In addition, our study introduces a different aspect of reliability that only becomes apparent when considering the full context of the use case (as we assessed users within the context of a specific task). When a dataset is modified, the overall model behaviour can be affected and must be considered. For TDA, this means that when modifying the dataset to addressing a particular error, users require explanations that not only suggest an effective modification but also ensure the model’s performance remains rather consistent.
@@
Our analysis underscores the importance of incorporating user perspectives to identify needs critical to the real-world application of this technology. Traditional TDA explanations only partially meet these needs, and further interdisciplinary efforts are required to develop methods with improved actionability.
@@ 5.2. Overlooked research topics for user-focused training data attribution
This work adds to a collection of human-centred explainable AI (HCXAI) works calling to centre XAI research around the user (e.g., (Brennen, 2020; Kaur et al., 2020; Rutjes et al., 2019; Ehsan et al., 2024; Mei et al., 2023; Lin et al., 2024)). By focusing on TDA explanations specifically, we identify several research areas that are necessary to study to advance TDA research towards a focus on users.
Mental models and TDA: Our study supports previous work (Rutjes et al., 2019; Kaur et al., 2020) that the most useful explanations depend on the user’s mental model of why the model errs: We find that what information is actionable to users depends in their hypothesis for reasons of model behaviour (Theme area: Action depends on user’s hypothesis). Therefore, future research should explore how TDA influences and is influenced by mental models, as well as how TDA can be adapted to reflect mental models (e.g. by defining relevance through the action of upweighting as opposed to removal), potentially leading to more aligned and actionable explanations that are well-defined.
User-centred group attribution: The survey findings reveal a clear need for attributing model behaviour to groups of training samples, as it provides more informative insights (Theme area: Group attribution preference). While existing approaches (Basu et al., 2020; Park et al., 2023; Koh et al., 2019) address this, our findings highlight the need for creating groupings that make sense to users (e.g., based on ground truth or predicted class), an aspect not yet widely explored in TDA research.
Holistic understanding of model errors: Our thematic analysis of survey responses highlights the importance of understanding the model error itself before attempting to fix it (Theme: Grouping erroneous test samples). If the root cause is a spurious correlation, the error likely extends beyond the current test sample and may not fully represent the issue. We recommend future research focus on understanding error types rather than isolated instances (e.g., slice discovery (Wang et al., 2024a)) and on attributing model behaviour on groups of test samples to training data (e.g., (Jain et al., 2024)).
Reliability of TDA: TDA quantifies the impact of training samples on model errors by predicting changes in model behaviour, serving as an explanation to the user. For the explanations to be informative, the attribution must be reliable. It is essential to understand how overall model behaviour will be affected as mentioned by our participants (Theme area: Reliability of explanations). We recommend future research to extend existing work on the fragility, reliability, and stability of TDA (Basu et al., 2021; Nguyen et al., 2023; Epifano et al., 2023; Bae et al., 2022) to address these needs.
@@ 5.3. Limitations
This study has several limitations. First, the use of purposive and snowball sampling techniques in participant recruitment may have introduced selection bias. Consequently, the sample may not be fully representative of the broader population of professionals working with machine learning or model developers in general. Second, the relatively small sample size in our studies limits the robustness of the quantitative analyses, so the results should be interpreted with caution and lead us to focus more strongly on the qualitative analysis. Third, we simulated the use of TDA explanations in a model debugging interface which are lab-like conditions for the survey study. Hence, the needfinding analysis is not rooted in the real workflows and is bound to our study conditions, including the interface design.
@@ 6. Conclusion
This paper proposes a top-down approach to conducting explainable AI (XAI) research which is driven by user needs as opposed to techno-centric solutions. We illustrate this approach with the emerging subfield of training data attribution (TDA) and present a needfinding study with AI practitioners. The study consists of two stages: First, we conduct semi-structured interviews with professionals who already work with AI systems (N=10) to gain an impression of common use cases and challenges. We find that TDA could especially benefit model developers, as their workflows are strongly data-centric. We derive a design space of TDA explanatory information that represents what relevant data for model debugging could look like. Second, we develop an interactive interface to extract user preferences in the design space and integrate it into a survey study (N=31) targeted at model developers. The mixed-methods analysis shows that the explanatory information needed from TDA for model debugging is highly dependent not only on the use case but also on the user and their intuitions. Finally, we highlight several research directions in user-focused TDA that are largely overlooked in the current landscape.
@@ Acknowledgements.
- Brennen (2020) Andrea Brennen. 2020. What Do People Really Want When They Say They Want ”Explainable AI?” We Asked 60 Stakeholders.. In Extended Abstracts of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI EA ’20). Association for Computing Machinery, New York, NY, USA, 1–7. https://doi.org/10.1145/3334480.3383047
@@
- Edwards (2005) Anthony WF Edwards. 2005. RA Fischer, statistical methods for research workers, (1925). In Landmark writings in western mathematics 1640-1940. Elsevier, 856–870.
@@
- Ehsan et al. (2024) Upol Ehsan, Samir Passi, Q. Vera Liao, Larry Chan, I-Hsiang Lee, Michael Muller, and Mark O Riedl. 2024. The Who in XAI: How AI Background Shapes Perceptions of AI Explanations. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 316, 32 pages. https://doi.org/10.1145/3613904.3642474
@@
- Guest et al. (2023) Greg Guest, Emily E. Namey, and Marilyn L. Mitchell. 2023. Collecting Qualitative Data: A Field Manual for Applied Research. SAGE Publications, Ltd, 55 City Road. https://doi.org/10.4135/9781506374680
- Guidotti et al. (2018) Riccardo Guidotti, Anna Monreale, Salvatore Ruggieri, Franco Turini, Fosca Giannotti, and Dino Pedreschi. 2018. A Survey of Methods for Explaining Black Box Models. ACM Comput. Surv. 51, 5, Article 93 (2018), 42 pages. https://doi.org/10.1145/3236009
- Guo et al. (2021) Han Guo, Nazneen Rajani, Peter Hase, Mohit Bansal, and Caiming Xiong. 2021. FastIF: Scalable Influence Functions for Efficient Model Interpretation and Debugging. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, Online and Punta Cana, Dominican Republic, 10333–10350. https://doi.org/10.18653/v1/2021.emnlp-main.808
- Hadash et al. (2022) Sophia Hadash, Martijn C. Willemsen, Chris Snijders, and Wijnand A. IJsselsteijn. 2022. Improving understandability of feature contributions in model-agnostic explainable AI tools. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 487, 9 pages. https://doi.org/10.1145/3491102.3517650
@@
- Hampel (1974) Frank R. Hampel. 1974. The Influence Curve and its Role in Robust Estimation. J. Amer. Statist. Assoc. 69, 346 (1974), 383–393. https://doi.org/10.1080/01621459.1974.10482962
@@
- Jin et al. (2024) Wei Jin, Haohan Wang, Daochen Zha, Qiaoyu Tan, Yao Ma, Sharon Li, and Su-In Lee. 2024. DCAI: Data-centric Artificial Intelligence. In Companion Proceedings of the ACM Web Conference 2024 (Singapore, Singapore) (WWW ’24). Association for Computing Machinery, New York, NY, USA, 1482–1485. https://doi.org/10.1145/3589335.3641297
- Juneja et al. (2024) Prerna Juneja, Wenjuan Zhang, Alison Marie Smith-Renner, Hemank Lamba, Joel Tetreault, and Alex Jaimes. 2024. Dissecting users’ needs for search result explanations. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 841, 17 pages. https://doi.org/10.1145/3613904.3642059
- Kaur et al. (2020) Harmanpreet Kaur, Harsha Nori, Samuel Jenkins, Rich Caruana, Hanna Wallach, and Jennifer Wortman Vaughan. 2020. Interpreting Interpretability: Understanding Data Scientists’ Use of Interpretability Tools for Machine Learning. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3313831.3376219
@@
- Kim et al. (2023) Sunnie S. Y. Kim, Elizabeth Anne Watkins, Olga Russakovsky, Ruth Fong, and Andrés Monroy-Hernández. 2023. ”Help Me Help the AI”: Understanding How Explainability Can Support Human-AI Interaction. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 250, 17 pages. https://doi.org/10.1145/3544548.3581001

### v2

@@ 5. Discussion
This work presents a needfinding study for training data attribution (TDA) explanations in model development to inform user-focused TDA research in a top-down approach. We summarise our findings to provide answers to the research questions (§ 5.1), identify research gaps that need to be addressed for more user-focused TDA (§ 5.2). We discuss the study framework in a general sense as an approach to bridge method and needfinding research especially in early development phases of a technology (§ 5.3), followed by limitations (§ 5.4).
@@ 5.1. Answers to the research questions
This study sets out to understand user needs for TDA explanations in model development. We pose three questions to understand use cases, explicit and latent needs, and understudied areas of research in user-focused TDA. We provide our answers based on our study below.
5.1.1. RQ1: What data-related explanation needs, both explicit and latent, do model developers have?
To address this question, we conducted a two-part user study comprising an interview study (§ 4.1) and a scenario-based interactive study (§ 4.4). We aimed to explore how data-related explanations are used in model development and to gain a deeper understanding of both explicit and latent needs.
Explanations are needed when the model behaves unexpectedly. Since ML development is largely data-driven, developers often inspect unexpected model behaviour from a data-centric perspective (Interview theme: Data-focused debugging). We observed a strong connection between a developer’s mental model of the ML system and their corresponding explanation needs. In practice, developers tend to form hypotheses, often data-related, to make sense of unexpected model behaviour (e.g., a developer may hypothesise that a model behaves unexpectedly due to data artefacts). Explanations serve as a means to confirm or reject these hypotheses and to refine or generate new ones, and inform subsequent actions taken by the developer to improve the system (Interactive study theme: Error hypotheses inform action).
Since each developer’s mental model is shaped by their experience and expectations, their explanation needs are also diverse. The needs TDA explanations must meet reflect developers’ data-centric information needs during model building. For example, developers seek insights to identify low-quality data when they suspect wrong labels to be the cause for a model error, or detect broader issues like distribution shifts when they believe this could have occured (Interview theme: Data quality), or recognise gaps where additional data is needed when they suspect their dataset to be insufficient for the model to learn the task (Interview theme: Data is most important).
Hence, these findings reveal that model developers have explicit needs for explanations that help identify possible causes of model error, whether by examining the impact of removing certain training samples or by highlighting the potential value of adding new data. At a deeper level, developers also express latent needs: explanations must be actionable, reliable, and aligned with their mental models. Explanations are not only used to understand the model’s behaviour, but also to guide meaningful decisions during development.
5.1.2. RQ2: To what extent do existing TDA approaches align with model developers’ needs?
To answer this question, we reflect on existing TDA approaches in light of the identified explicit and latent needs: TDA explanations need to be actionable, reliable, and aligned with the developer’s mental models. From these needs, we identify different TDA tasks that developers need TDA explanations to fulfil, next to the overarching need for reliability (see Table 4).
@@ 5.1.2. RQ2: To what extent do existing TDA approaches align with model developers’ needs?
| Influence attribution (e.g., (Koh and Liang, 2017; Pruthi et al., 2020; Guo et al., 2021; Schioppa et al., 2022; Grosse et al., 2023; Park et al., 2023; Bae et al., 2024)) | Removal | Sample loss | 1 | 1 | 0 |
@@
Existing TDA approaches partly address explicit needs.
The formal definition of TDA (§ 3) characterises attribution as the change in model output resulting from training with versus without a specific training sample (Hampel, 1974). (Koh and Liang, 2017; Grosse et al., 2023; Park et al., 2023; Bae et al., 2024; Choe et al., 2024; Kwon et al., 2023; Pruthi et al., 2020) focus on efficiently and faithfully estimating this quantity (influence attribution in Table 4) or its extension to groups of training samples (Basu et al., 2020; Park et al., 2023; Koh et al., 2019) (group influence attribution in Table 4). TDA has been applied to a variety of tasks, including the identification of mislabeled data, analysis of adversarial vulnerabilities, and detection of domain mismatches (Koh and Liang, 2017). Mislabel detection, in particular, is a widely used evaluation across several studies (Pruthi et al., 2020; Park et al., 2023; Schioppa et al., 2022), and TDA methods have demonstrated the ability to identify outliers and mislabeled samples in training data (Koh and Liang, 2017; Pruthi et al., 2020). These capabilities align with certain explicit developer needs, particularly those focused on identifying low-quality data to remove. However, since TDA quantifies a sample’s relevance through the lens of its removal, it is especially well-suited for scenarios involving data elimination, such as detecting harmful or noisy samples. In contrast, broader data-centric explanation needs, such as identifying distribution shifts or detecting underrepresented regions in the data, are generally not addressed by current TDA methods, as they are not typically designed with these needs in mind.
Existing TDA work does not consider developers’ latent need for actionable information.
Our study shows that developers have the latent need for data-based explanations to enable them to confirm or revise their mental models of model behaviour, which are diverse. Fulfilling this latent need is tied to the developer’s mental model and makes an explanation actionable. This human-centred aspect of actionability is rarely considered in existing TDA work. While existing TDA methods are useful for measuring the impact of removing and retraining on individual samples, because they are studied to quantify this objective, developers may require different forms of attribution to address a wider range of questions during model development (e.g., lacking data attribution in Table 4).
Existing TDA work partly implicitly addresses developers’ need for reliability.
@@
Regarding the need for reliable information, existing TDA methods have been criticised for their low reliability in deep learning (Nguyen et al., 2023; Basu et al., 2021; Bae et al., 2022). In addition, our study introduces a different aspect of reliability that only becomes apparent when considering the full context of the use case (as we assessed participants within the context of a specific task). When a dataset is modified, the overall model behaviour can be affected and must be considered. For TDA, this means that when modifying the dataset to address a particular error, developers require explanations that not only suggest an effective modification but also ensure the model’s performance remains rather consistent.
@@
Our analysis underscores the importance of incorporating user perspectives to identify needs critical to the real-world application of this technology. Traditional TDA explanations only partially meet these needs, and further interdisciplinary efforts beyond method research are required to develop methods with improved actionability.
5.1.3. RQ3: What kind of information produced by TDA will be useful for model developers?
To answer this question, we examine the gap between the needs we found and the needs that existing TDA work meets (see Table 4).
Our study reveals that developers need different types of information to understand model behaviour, depending on their mental models and hypotheses of why the model errs during development, in line with (Cabrera et al., 2023). This necessitates TDA explanations that are flexible and adaptable to individual needs. These needs vary based on the application domain and the developer’s mental models that are shaped by the developer’s experiences and expectations. Experienced developers may form hypotheses about model errors based on their past experiences. In contrast, less experienced developers may form hypotheses based on what they learnt previously about machine learning models. TDA that aligns with the developer’s mental models will be useful.
Developers prefer group attribution, as groups provide more information than individual samples and are expected to paint a more conclusive picture of the model behaviour. Hence, information related to group attribution will be useful for model developers.
Additionally, developers need TDA explanations to be reliable, considering how data-centric actions might affect the model as a whole. Reliability entails two main aspects: First, the information needs to be reliable for it to be actionable and effective for the task (i.e., model debugging). Second, TDA explanations should also account for any broader changes in the model’s behaviour that may affect the attribution to training samples.
@@ 5.2. Overlooked research topics for user-focused training data attribution
This work adds to a collection of human-centred explainable AI (HCXAI) works calling to centre XAI research around users (e.g., (Brennen, 2020; Kaur et al., 2020; Rutjes et al., 2019; Ehsan et al., 2024; Mei et al., 2023; Lin et al., 2024)). We focus particularly on model developer users, as they have agency over the dataset. By focusing on TDA explanations specifically for model developers, we identify several research areas that are necessary to study to advance TDA research towards a focus on developers.
Mental models and TDA: Our study supports previous work (Rutjes et al., 2019; Kaur et al., 2020) that the most useful explanations depend on the user’s mental model of why the model errs: We find that what information is actionable to users depends in their hypothesis for reasons of model behaviour (Theme: Error hypothesis informs action). Therefore, future research should explore how TDA influences and is influenced by mental models, as well as how TDA can be adapted to reflect mental models (e.g., by defining relevance through the action of upweighting as opposed to removal), potentially leading to more aligned and actionable explanations that are well-defined.
User-centred group attribution: The survey findings reveal a clear need for attributing model behaviour to groups of training samples, as it provides more informative insights (Theme: Group for global effect). While existing approaches (Basu et al., 2020; Park et al., 2023; Koh et al., 2019) address this, our findings highlight the need for creating groupings that make sense to users (e.g., based on ground truth or predicted class), an aspect not yet widely explored in TDA research.
Holistic understanding of model errors: Our thematic analysis of survey responses highlights the importance of understanding the model error itself before attempting to fix it (Theme: Inspection actions, Group precision). If the root cause is a spurious correlation, the error likely extends beyond the current test sample and may not fully represent the issue. We recommend future research focus on understanding error types rather than isolated instances (e.g., slice discovery (Wang et al., 2024)) and on attributing model behaviour on groups of test samples to training data (e.g., (Jain et al., 2024)).
Reliability of TDA: TDA quantifies the impact of training samples on model errors by predicting changes in model behaviour, serving as an explanation to the user. For the explanations to be informative, the attribution must be reliable. It is essential to understand how overall model behaviour will be affected, as mentioned by our participants. We recommend future research to extend existing work on the fragility, reliability, and stability of TDA (Basu et al., 2021; Nguyen et al., 2023; Epifano et al., 2023; Bae et al., 2022) to address these needs.
@@
5.3. A framework for needs-based research
In this work, we demonstrate a three-step process to identify potentially overlooked research directions and inform needs-based research for socially situated technology like machine learning model explanations:
Understanding the user: The first step aims to identify the potential user of a technology: Examining who may have an interest in the technology under consideration of their agency, as well as what purpose the technology serves. In this work, we consider XAI technology. Hence, potential users interact with machine learning systems and have an interest in understanding them. Often in XAI, the users of ML technology in high-risk application areas are selected as the main audience due to legal requirements of transparency by the EU AI Act (Parliament, 2023) for example, but interested users could extend beyond this group to any interested users. Since we specifically focus on TDA explanations, we were able to narrow the user group to model developers who have agency over the training data.
Understanding their needs: The second step aims to understand the needs of users through qualitative methods like semi-structured interviews with thematic coding. By engaging with users, we can understand the application context of the technology, including existing workflows that the technology needs to fit in and existing challenges that the technology could potentially help with. Especially in the case of novel technology that is not yet used, it is important to identify potential use cases to understand what needs exist. In this work, we were able to analyse common use cases of TDA explanations in model development from interviews with machine learning developers (§ 4.1) and derive the design space of TDA explanations (§ 4.3).
Validate their needs: The third step aims to validate the needs found in the previous step through a prototype or scenario-based study, so that users can experience the technology. This step enables analysis that is rooted in real interaction, and shall provide deeper insights and validation of the user needs. In this work, we designed a mock-up interface for the scenario-based interactive user study and were able to deepen our understanding of what kind of information is needed from TDA explanations (§ 4.4). We were able to clearly identify preferences for group attribution, as well as currently understudied research directions, such as the need for explanations of model behaviour that point to missing data.
This process is intended to build a bridge between user research and methods research, as it allows for potential users to understand the capabilities of novel methods, and results in open questions and example use cases for method researchers. While this paper applies the framework to XAI, in particular TDA explanations, it is a general approach that could be applied to any technology intended for use in a sociotechnical context.
@@ 5.4. Limitations
This study has several limitations. First, the use of purposive and snowball sampling techniques in participant recruitment may have introduced selection bias. Consequently, the sample may not be fully representative of the broader population of professionals working with machine learning or model developers in general. Second, the relatively small sample size in our studies limits the robustness of the quantitative analyses, so the results should be interpreted with caution and lead us to focus more strongly on the qualitative analysis. Third, we simulated the use of TDA explanations in a model debugging interface, which are lab-like conditions for the survey study. Hence, the needfinding analysis is not rooted in the real workflows and is bound to our study conditions, including the interface design.
Yet, we remark that this third limitation is somewhat systematic, as TDA explanations are not widely adopted in real workflows yet. At the time of the studies, our participants were not familiar with TDA explanations at all, and the first libraries for data attribution were only recently published (Bareeva et al., 2024; Deng et al., 2024). We therefore believe that future studies will overcome this limitation as more real-world users engage with the TDA pipeline in their work.
@@ 6. Conclusion
This paper proposes a top-down approach to conducting explainable AI (XAI) research which is driven by user needs as opposed to techno-centric solutions. We illustrate this approach with the emerging subfield of training data attribution (TDA) and present a needfinding study with AI practitioners. The study consists of two stages: First, we conduct semi-structured interviews with machine learning developers in high-risk domain areas (N=6) to gain an understanding of common use cases and challenges. We derive a design space of TDA explanatory information that represents what relevant data for model debugging could look like. Second, we develop an interactive interface to explore developer preferences in the design space through a scenario-based interactive user study (N=31) targeted at model developers. The second study allows us to validate and expand the needfinding analysis. The mixed-methods analysis shows that the explanatory information needed from TDA for model debugging is highly dependent not only on the use case but also on the developer and their mental model and intuitions. Finally, we establish our study approach as a framework for needs-based research and highlight several research directions in user-focused TDA that are largely overlooked in the current landscape.
@@ Acknowledgements.
- Agarwal et al. (2017) Naman Agarwal, Brian Bullins, and Elad Hazan. 2017. Second-order stochastic optimization for machine learning in linear time. Journal of Machine Learning Research 18, 116 (2017), 1–40.
@@
- Arnoldi (1951) Walter Edwin Arnoldi. 1951. The principle of minimized iterations in the solution of the matrix eigenvalue problem. Quarterly of applied mathematics 9, 1 (1951), 17–29.
@@
- Bareeva et al. (2024) Dilyara Bareeva, Galip Ümit Yolcu, Anna Hedström, Niklas Schmolenski, Thomas Wiegand, Wojciech Samek, and Sebastian Lapuschkin. 2024. Quanda: An Interpretability Toolkit for Training Data Attribution Evaluation and Beyond. arXiv:2410.07158 [cs.LG] https://arxiv.org/abs/2410.07158
@@
- Brennen (2020) Andrea Brennen. 2020. What Do People Really Want When They Say They Want ”Explainable AI?” We Asked 60 Stakeholders.. In Extended Abstracts of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI EA ’20). Association for Computing Machinery, New York, NY, USA, 1–7. doi:10.1145/3334480.3383047
@@
- Cabrera et al. (2023) Ángel Alexander Cabrera, Marco Tulio Ribeiro, Bongshin Lee, Robert Deline, Adam Perer, and Steven M. Drucker. 2023. What Did My AI Learn? How Data Scientists Make Sense of Model Behavior. ACM Trans. Comput.-Hum. Interact. 30, 1, Article 1 (March 2023), 27 pages. doi:10.1145/3542921
@@
- Deng et al. (2024) Junwei Deng, Ting-Wei Li, Shiyuan Zhang, Shixuan Liu, Yijun Pan, Hao Huang, Xinhe Wang, Pingbang Hu, Xingjian Zhang, and Jiaqi Ma. 2024. dattri: A Library for Efficient Data Attribution. Advances in Neural Information Processing Systems 37 (2024), 136763–136781.
@@
- Ehsan et al. (2024) Upol Ehsan, Samir Passi, Q. Vera Liao, Larry Chan, I-Hsiang Lee, Michael Muller, and Mark O Riedl. 2024. The Who in XAI: How AI Background Shapes Perceptions of AI Explanations. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 316, 32 pages. doi:10.1145/3613904.3642474
@@
- Engstrom et al. (2025) Logan Engstrom, Andrew Ilyas, Benjamin Chen, Axel Feldmann, William Moses, and Aleksander Madry. 2025. Optimizing ml training with metagradient descent. arXiv preprint arXiv:2503.13751 (2025).
@@
- Ghorbani and Zou (2019) Amirata Ghorbani and James Zou. 2019. Data shapley: Equitable valuation of data for machine learning. In International conference on machine learning. PMLR, 2242–2251.
@@
- Guest et al. (2023) Greg Guest, Emily E. Namey, and Marilyn L. Mitchell. 2023. Collecting Qualitative Data: A Field Manual for Applied Research. SAGE Publications, Ltd, 55 City Road. doi:10.4135/9781506374680
- Guidotti et al. (2018) Riccardo Guidotti, Anna Monreale, Salvatore Ruggieri, Franco Turini, Fosca Giannotti, and Dino Pedreschi. 2018. A Survey of Methods for Explaining Black Box Models. ACM Comput. Surv. 51, 5, Article 93 (2018), 42 pages. doi:10.1145/3236009
- Guo et al. (2021) Han Guo, Nazneen Rajani, Peter Hase, Mohit Bansal, and Caiming Xiong. 2021. FastIF: Scalable Influence Functions for Efficient Model Interpretation and Debugging. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, Online and Punta Cana, Dominican Republic, 10333–10350. doi:10.18653/v1/2021.emnlp-main.808
- Hadash et al. (2022) Sophia Hadash, Martijn C. Willemsen, Chris Snijders, and Wijnand A. IJsselsteijn. 2022. Improving understandability of feature contributions in model-agnostic explainable AI tools. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 487, 9 pages. doi:10.1145/3491102.3517650
@@
- Hampel (1974) Frank R. Hampel. 1974. The Influence Curve and its Role in Robust Estimation. J. Amer. Statist. Assoc. 69, 346 (1974), 383–393. doi:10.1080/01621459.1974.10482962
- Hara et al. (2019) Satoshi Hara, Atsushi Nitanda, and Takanori Maehara. 2019. Data cleansing for models trained with SGD. Advances in Neural Information Processing Systems 32 (2019).
@@
- Hohman et al. (2020) Fred Hohman, Kanit Wongsuphasawat, Mary Beth Kery, and Kayur Patel. 2020. Understanding and Visualizing Data Iteration in Machine Learning. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–13. doi:10.1145/3313831.3376177
- Hong et al. (2020) Sungsoo Ray Hong, Jessica Hullman, and Enrico Bertini. 2020. Human Factors in Model Interpretability: Industry Practices, Challenges, and Needs. Proc. ACM Hum.-Comput. Interact. 4, CSCW1, Article 68 (May 2020), 26 pages. doi:10.1145/3392878
- Hu et al. (2024) Yuzheng Hu, Pingbang Hu, Han Zhao, and Jiaqi Ma. 2024. Most Influential Subset Selection: Challenges, Promises, and Beyond. In Advances in Neural Information Processing Systems, A. Globerson, L. Mackey, D. Belgrave, A. Fan, U. Paquet, J. Tomczak, and C. Zhang (Eds.), Vol. 37. Curran Associates, Inc., 119778–119810. https://proceedings.neurips.cc/paper_files/paper/2024/file/d8684e49752e06ac5e4b554b60ad212a-Paper-Conference.pdf
- Ilyas and Engstrom (2025) Andrew Ilyas and Logan Engstrom. 2025. MAGIC: Near-Optimal Data Attribution for Deep Learning. arXiv preprint arXiv:2504.16430 (2025).
@@
- Jin et al. (2024) Wei Jin, Haohan Wang, Daochen Zha, Qiaoyu Tan, Yao Ma, Sharon Li, and Su-In Lee. 2024. DCAI: Data-centric Artificial Intelligence. In Companion Proceedings of the ACM Web Conference 2024 (Singapore, Singapore) (WWW ’24). Association for Computing Machinery, New York, NY, USA, 1482–1485. doi:10.1145/3589335.3641297
- Juneja et al. (2024) Prerna Juneja, Wenjuan Zhang, Alison Marie Smith-Renner, Hemank Lamba, Joel Tetreault, and Alex Jaimes. 2024. Dissecting users’ needs for search result explanations. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 841, 17 pages. doi:10.1145/3613904.3642059
- Kaur et al. (2020) Harmanpreet Kaur, Harsha Nori, Samuel Jenkins, Rich Caruana, Hanna Wallach, and Jennifer Wortman Vaughan. 2020. Interpreting Interpretability: Understanding Data Scientists’ Use of Interpretability Tools for Machine Learning. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–14. doi:10.1145/3313831.3376219
@@
- Kim et al. (2023) Sunnie S. Y. Kim, Elizabeth Anne Watkins, Olga Russakovsky, Ruth Fong, and Andrés Monroy-Hernández. 2023. ”Help Me Help the AI”: Understanding How Explainability Can Support Human-AI Interaction. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 250, 17 pages. doi:10.1145/3544548.3581001
@@
- Krippendorff (2011) Klaus Krippendorff. 2011. Computing Krippendorff’s alpha-reliability.